Joe slpit his $20000 bonus between two mutual funds. The ace fund earned a 6% profit during the first year,while the loser fund lost 5%. The combined result of both investments was a $340 loss. How much did Joe invest in each fund.
Answer by checkley77(12844)

.06x-.05(20,000-x)=-340
.06x-1,000+.05x=-340
.11x+.05x=-340+1,000
.11x=660
x=660/.11
x=6,000 amount invested @ 6%.
20,000-6,000=14,000 invested @ 5%.
Proof:
.06*6,000-.05(14,0600)=-350
360-700=-340
-340=-340